# How to Learn a Star: Binary Classification with Starshaped Polyhedral Sets

### v1

@@ A.1 Examples of polyhedral fans (Section 2)
Figure 3: The functions $[\mathbf{x}]^{\diamond}$ and $[\mathbf{x}]^{B}$ restricted to the full-dimensional cones of the $2$-dimensional fans from Sections A.1 and A.1. $\mathbf{0}_{k}$ denotes the $k$-dimensional 00-vector.
[…]
@@
The vector $[\mathbf{x}]^{B}$ can be recovered from (4). The $2$-dimensional fan $B$ and the function $[\mathbf{x}]^{B}$ restricted to each maximal cone is depicted in Figure 3(b).
A.2 Examples of the geometry of the parameter space (Section 3)
[…]
@@ A.2 Examples of the geometry of the parameter space (Section 3)
$[-z]^{\Delta}=\left(\begin{smallmatrix}z\\ 0\end{smallmatrix}\right)\quad\text{ and }\quad[z]^{\Delta}=\left(\begin{smallmatrix}0\\ z\end{smallmatrix}\right).$
The associated data arrangement is depicted in Figure 4(a), and Figure 4(b) shows the $1$-dimensional dataset together with the stars associated to the points
@@
||||$\mathbf{x}^{(1)}$$\mathbf{x}^{(2)}$$\mathbf{x}^{(3)}$$\mathbf{x}^{(4)}$$\mathbf{x}^{(5)}$$\mathbf{x}^{(6)}$$\mathbf{x}^{(7)}$$\mathbf{x}^{(8)}$
@@
Figure 4: The data arrangement and dataset from Section A.2.
By Section 3.1, each half-open chamber of the data arrangement is the set of classification vectors $\mathbf{a}$ whose induced classifiers agree on the dataset. Thus, the number of false positives $\operatorname{FP}(\mathbf{a})$, the number of false negatives $\operatorname{FN}(\mathbf{a})$ and the discrete loss function $\operatorname{err}(\mathbf{a})$ are constant on each of the half-open chambers (cf. Section 3.1). Figure 5 shows the values of these functions and it can be verified that the sublevel sets of $\operatorname{FP}(\mathbf{a})$ and $\operatorname{FN}(\mathbf{a})$ are star-convex with centers $\mathbf{0}$ and $\infty$, respectively (cf. Theorem 3.2). As the data is perfectly separable, Theorem 3.3 implies that the sublevel sets of $\operatorname{err}(\mathbf{a})$ are star-convex and connected through walls of codimension $1$, as depicted in Figure 5(c).
@@
Figure 5: The values of $\operatorname{FP}(\mathbf{a}),\operatorname{FN}(\mathbf{a})$ and $\operatorname{err}(\mathbf{a})$ on the half-open chambers of the data arrangement for the $1$-dimensional dataset from Section A.2. The diagonal lines in all three images show the function $\lambda\mapsto\mathbf{a}^{*}(\lambda)$ defined in Section 3.2.
Figure 6 shows the level sets of the exponential loss function for the same example, for two choices of the rate parameter $\lambda$. In accordance to Section 3.2, the plots illustrate that the superlevel sets of these functions are convex. Moreover, the rate parameter has an influence on the (unique) maximizer of these functions, but they lie on a line $\lambda\mapsto\mathbf{a}^{*}(\lambda)$, as shown in Theorem 3.7.
The same line is drawn in Figures 5(a) and 5(b), illustrating that the numbers of false positives and false negative are monotone increasing and decreasing, respectively (cf. Section 3.2). In Figure 5(c), the line crosses regions where the $0/1$-loss has values $3,2,3,2,3,2,3,4$, respectively. This shows that the $0/1$-loss is not “convex” in the sense that the sequence of values of the $0/1$-loss along this line is unimodal, but goes up and down repeatedly.
@@
Figure 6: Level sets of exponential loss functions for different choices of $\lambda$ on the dataset from Section A.2. The black dot depicts the minimum $\mathbf{a}^{*}(\lambda)$, lying on the line of maxima when varying the choice of $\lambda$.
A.3 Examples of the geometry of the generalized parameter space (Section 4)
@@
Figure 7: The data points from Section A.3, with fixed $\mathbf{a}$ and two perfect classifiers.
[…]
Consider the $2$-dimensional labeled dataset $X$ consisting of $3$ distinct points $(\mathbf{x}^{(1)},y^{(1)})=(1,1,0),\ (\mathbf{x}^{(2)},y^{(2)})=(2,2,1),\ (\mathbf{x}^{(3)},y^{(3)})=(3,3,0),$ and let $\Delta$ be the $2$-dimensional Coxeter fan of type B (cf. Section A.1), which has $8$ rays. We fix the star through values $\mathbf{a}=\left(\frac{1}{3},3,\frac{1}{3},3,\frac{1}{3},3,\frac{1}{3},3\right).$
Figure 7 shows two different translations $\operatorname{star}(\mathbf{a})+\mathbf{t}$ that are perfect classifiers. Figure 8 shows the associated star arrangement in $\mathbb{R}^{2}$, and the sublevel sets of the translational 0/1-loss. In particular, in contrast to Theorem 3.3, it can be observed that the $0^{\text{th}}$ level set $L(\operatorname{err}_{\mathbf{a}},0)$ consists of two full-dimensional connected components. The translated stars in Figure 7 show one example from each of the connected components.
222233333333333333333333333322222222222222222222222222222222111111111111111111110000-2-2-1-100112233445566
Figure 8: The star arrangement from Section A.3, and the level sets of the translational 0/1-loss function.
Appendix B Proofs
B.1 Proof of Theorem 3.1
@@ Appendix B Proofs
Then $f_{\mathbf{a}}^{\Delta}(\mathbf{v}_{i})=a_{i}\geq 1$ if and only if $\ell(\mathbf{v}_{i})=1$, and thus $c_{\mathbf{a}}(\mathbf{v}_{i})=\ell(\mathbf{v}_{i})$ as claimed. It follows that the set of classifiers shatters the set of generators and thus the VC-dimension of $S^{\Delta}$ is at least $n$.
To see that the VC dimension is at most $n$, we assume that there is a set $\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n+1)}\in\mathbb{R}^{d}$ of $n+1$ points that can be shattered by $S^{\Delta}$. By construction, for all $11\leq i\leq n+$ 1, $c_{\mathbf{a}}(\mathbf{x}^{(i)})=0$ if and only if $\langle[\mathbf{x}^{(i)}]^{\Delta},\mathbf{a}\rangle\leq 1$. In particular, if $\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n+1)}$ can be shattered by $S^{\Delta}$ then $[\mathbf{x}^{(1)}]^{\Delta},\ldots,[\mathbf{x}^{(n+1)}]^{\Delta}$ can be shattered by the set of half-spaces of the form $\{\mathbf{x}\in\mathbb{R}^{n}\colon\mathbf{a}^{T}\mathbf{x}\leq b,a_{1},\ldots,a_{n}\geq 0\}$. This is a subset of the set of halfspaces considered in Section B.1 below, and thus we obtain a contradiction by Section B.1. This completes the proof. ∎
[…]
@@
Then the VC-dimension of $\mathcal{H}$ equals $n$.
@@
We need to show that there is no set of $n+1$ that can be shattered by $\mathcal{H}$. Let $\mathbf{y}_{1},\ldots,\mathbf{y}_{n+1}$ be points in $\mathbf{R}^{n}$ and let $\tilde{\mathbf{y}}_{1},\ldots,\tilde{\mathbf{y}}_{n+1}$ be their projections on the first $n-1$ coordinates. Then $\tilde{\mathbf{y}}_{1},\ldots,\tilde{\mathbf{y}}_{n+1}$ are affinely dependent, that is, there exist $\mu_{1},\ldots,\mu_{n}$, not all equal to zero, such that $\sum_{i}\mu_{i}=0$ and $\sum_{i}\mu_{i}\tilde{\mathbf{y}}_{i}=\mathbf{0}$. Let
$A\ =\ \sum_{i\colon\mu_{i}>0}\mu_{i}\ =\ \sum_{i\colon\mu_{i}<0}(-\mu_{i})\,.$
@@
Then both $\mathbf{w}_{+}$ and $\mathbf{w}_{-}$ agree with $\tilde{\mathbf{w}}$ on the first $n-1$ coordinates. W.l.o.g we may assume that $(w_{-})_{n}\leq(w_{+})_{n}$, that is, the last coordinate of $w_{-}$ is not bigger than the last coordinate of $\mathbf{w}_{+}$. Then we claim that there exists no half-space $\{\mathbf{x}\in\mathbb{R}^{n}\colon\mathbf{a}^{T}\mathbf{x}\leq b,a_{n}\geq 0\}$ in $\mathcal{H}$ such that $\mathbf{a}^{T}\mathbf{y}_{i}\leq b$ for all $i$ such that $\mu_{i}>0$ and $\mathbf{a}^{T}\mathbf{y}_{i}>b$ for all $i$ such that $\mu_{i}<0$. To see this, we assume to the contrary that such a hyperplane exists. We then have
@@
and similarly $\mathbf{a}^{T}\mathbf{w}_{-}>b$. In particular, $\mathbf{a}^{T}\mathbf{w}_{-}>\mathbf{a}^{T}\mathbf{w}_{+}$. Since $\mathbf{w}_{+}$ and $\mathbf{w}_{-}$ agree on the first $n-1$ coordinates it therefore follows that $a_{n}(w_{+})_{n}<a_{n}(w_{-})$ and thus $\mathbf{w}_{+}<\mathbf{w}_{-}$ since $a_{n}>0$, a contradiction. Thus, no hyperplane in $\mathcal{H}$ satisfies the claim and thus $\mathcal{H}$ does not shatter any set of $n+1$ points. The VC-dimension is thus at most $n$. To see that it is fact equal to $n$ we observe that the set of unit vectors $\mathbf{e}_{1},\ldots,\mathbf{e}_{n}$ can be shattered. ∎
B.2 Proof of Theorem 3.2
@@
From the proof of Theorem 3.4 we see that the Hessian of $\mathcal{L}(\mathbf{a})$ is a negative linear combination of the rank-$1$ matrices $\left([\mathbf{x}^{(i)}]^{\Delta}\right)^{T}[\mathbf{x}^{(i)}]^{\Delta}$ for $\mathbf{x}^{(i)}\in X_{1}$, that is, $\operatorname{Hess}\mathcal{L}(\mathbf{a})=\sum_{i}\lambda_{i}\left([\mathbf{x}^{(i)}]^{\Delta}\right)^{T}[\mathbf{x}^{(i)}]^{\Delta}$ for some $\lambda_{i}<0$ where the sum is over all $i$ such that $\mathbf{x}^{(i)}\in X_{1}$. Now let $\mathbf{v}$ be an eigenvector of $\operatorname{Hess}\mathcal{L}(A)$ with eigenvalue $\mu$. Since $A_{X_{1}}$ has rank $n$, there exists an $\mathbf{x}^{(i_{0})}\in X_{1}$ with $\langle[\mathbf{x}^{(i_{0})}]^{T},\mathbf{v}\rangle\neq 0$. Then
@@
It follows that $\mu\neq 0$ and thus $\mu<0$. Thus, $\operatorname{Hess}\mathcal{L}(A)$ is negative definite and thus $\mathcal{L}(A)$ is strictly concave. Since the parameter space $\mathbb{R}_{>0}^{n}=\left\{\mathbf{a}:\mathbf{a}>0\right\}$ is convex, $\operatorname{Hess}\mathcal{L}(A)$ has a unique maximum on the extended positive orthant $(\mathbb{R}_{\geq 0}\cup\{\infty\})^{n}$. If furthermore the rank of $A_{X_{0}}$ is $n$ then for each $j\in\{1,\dots,n\}$ there exists an $x^{(i)}\in X_{0}$ such that $[\mathbf{x}^{(i)}]_{j}>0$. Therefore, we see that $\mathcal{L}(\mathbf{a})\rightarrow-\infty$ whenever $a_{j}\rightarrow\infty$. Thus, the unique maximum must be attained in the closed positive orthant $(\mathbb{R}_{\geq 0})^{n}$. ∎
B.6 Proof of Theorem 3.7
@@
From the proof of Theorem 3.7 we see that for any $0<\lambda<\lambda^{\prime}$ holds $\mathbf{a}^{*}(\lambda^{\prime})=\lambda/\lambda^{\prime}\cdot\mathbf{a}^{*}(\lambda)<\mathbf{a}^{*}(\lambda)$. Since $\mathbf{a}^{*}(\lambda)\in\mathbb{R}_{>0}^{n}$ both claims follow from Theorem 3.2. ∎
B.8 Proof of Section 4.1
@@
We begin with the first statement. First, let $C=\operatorname{cone}(\mathbf{v}_{i_{1}},\dots,\mathbf{v}_{i_{d}})$ be a full-dimensional cone of $\Delta$ such that $\mathbf{t}$ is contained in the interior of the cone $\mathbf{x}^{(i)}-C$ of the fan $\mathbf{x}^{(i)}-\Delta$. Equivalently, $\mathbf{x}^{(i)}-\mathbf{t}\in C$. With $V_{C}=(\mathbf{v}_{i_{1}}\dots\mathbf{v}_{i_{d}})$, we can thus compute $[\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta}=V_{C}^{-1}(\mathbf{x}^{(i)}-\mathbf{t})$, which is a linear function in $\mathbf{t}$. The statement for lower-dimensional cones follows by taking limits. The second statement follows from substitution of the variables $\mathbf{x}\mapsto-\mathbf{t}$ and $\mathbf{t}\mapsto\mathbf{x}^{(i)}$ in (3). ∎
B.9 Proof of Theorem 4.1
@@
Section 4.1 implies that
@@
By Section 4.1, for any $i$, $\mathbf{t}\mapsto f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)}-\mathbf{t})=\langle\mathbf{a},[\mathbf{x}^{(i)}-\mathbf{t}]\rangle$ is linear in $\mathbf{t}$ on every maximal cell $F$ of the fan arrangement of $\mathbf{x}^{(i)}-\Delta$, $i=1,\ldots,m$. Let $g_{F}(\mathbf{t})=\langle\mathbf{a},[\mathbf{x}^{(i)}-\mathbf{t}]\rangle$ be this linear function. Then the Hessian of $\mathcal{L}_{\mathbf{a}}(\mathbf{t})$ is the sum of matrices of the form
@@
The first set equals >$\mathbb{R}^{n}_{>0}\times(-C+\mathbf{x}^{(i)})$. Restricted to $(\mathbf{a},\mathbf{t})$ such that $\mathbf{t}\in\mathbf{x}^{(i)}-C$, the function $\mathbf{t}\mapsto[\mathbf{x}^{(i)}-\mathbf{t}]$ is a linear map by Section 4.1, so the expression $f^{\Delta}_{\mathbf{a}}(\mathbf{x}^{(i)}-\mathbf{t})=\langle[\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta},\mathbf{a}\rangle$ is a quadratic polynomial in variables $t_{1},\dots,t_{d},a_{1},\dots,a_{n}$. Therefore, $\mathcal{S}^{0}(C,\mathbf{x}^{(i)})$ it the intersection of solutions to the linear inequalities defining the polyhedral cone >$\mathbb{R}^{n}_{>0}\times(-C+\mathbf{x}^{(i)})$, and the quadratic inequality $\langle[\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta},\mathbf{a}\rangle\leq 1$. Similarly, $\mathcal{S}^{1}(C,\mathbf{x}^{(i)})$ is the intersection of >$\mathbb{R}^{n}_{>0}\times(-C+\mathbf{x}^{(i)})$ with the set of solutions to the inequality $\langle[\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta},\mathbf{a}\rangle>1$. ∎
B.12 Proof of Theorem 4.3
@@
We consider the subdivision of >$\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$ into sets
@@
where we range over all possible $b(C,i)\in\{0,1\}$ for all $C\in\Delta,i\in\{1,\dots,m\}$. By Section 4.2, each $S^{b(C,i)}(C,\mathbf{x}^{(i)})$ is basic semialgebraic with defining polynomials of degree at most $2$, and hence the same holds for the above finite intersection. By construction, the extended $0/1$-loss is constant on each $\bigcap_{C\in\Delta}\ \bigcap_{i=1}^{m}\mathcal{S}^{b(C,i)}(C,\mathbf{x}^{(i)})$. Fix $k\in\mathbb{Z}_{\geq 0}$. Then the $k^{\text{th}}$ level set $L(\operatorname{err},k)$ of the extended $0/1$-loss is the (finite) union over all sets $\bigcap_{C\in\Delta}\ \bigcap_{i=1}^{m}\mathcal{S}^{b(C,i)}(C,\mathbf{x}^{(i)})$ on which the extended $0/1$-loss is equal to $k$. Thus, the level set is a semialgebraic set. Since sublevel sets are finite unions of level sets, the same holds for sublevel sets. ∎
B.13 Proof of Theorem 4.4

### v2

@@ A.1 Examples of polyhedral fans (Section˜2)
Figure 4: The functions $\mathbf{x}]^{\diamond}$ and $\mathbf{x}]^{B}$ restricted to the full-dimensional cones of the $2$-dimensional fans from Sections˜A.1 and A.1. $\mathbf{0}_{k}$ denotes the $k$-dimensional 0-vector.
Example A.2 (Type B stars).
@@
The vector $\mathbf{x}]^{B}$ can be recovered from (4). The $2$-dimensional fan $B$ and the function $\mathbf{x}]^{B}$ restricted to each maximal cone is depicted in Figure˜4(b).
A.2 Examples of the geometry of the parameter space (Section˜3)
Example A.3 (Classifications of 1-dimensional dataset).
@@ A.2 Examples of the geometry of the parameter space (Section˜3)
[…]
The associated data arrangement is depicted in Figure˜5(a), and Figure˜5(b) shows the $1$-dimensional dataset together with the stars associated to the points
@@
[…]
@@
Figure 5: The data arrangement and dataset from Section˜A.2.
By Section˜3.1, each half-open chamber of the data arrangement is the set of classification vectors $\mathbf{a}$ whose induced classifiers agree on the dataset. Thus, the number of false positives $\operatorname{FP}(\mathbf{a})$, the number of false negatives $\operatorname{FN}(\mathbf{a})$ and the discrete loss function $\operatorname{err}(\mathbf{a})$ are constant on each of the half-open chambers (cf. Section˜3.1). Figure˜6 shows the values of these functions and it can be verified that the sublevel sets of $\operatorname{FP}(\mathbf{a})$ and $\operatorname{FN}(\mathbf{a})$ are star-convex with centers $\mathbf{0}$ and $\infty$, respectively (cf. Theorem˜3.4). As the data is perfectly separable, Theorem˜3.5 implies that the sublevel sets of $\operatorname{err}(\mathbf{a})$ are star-convex and connected through walls of codimension $1$, as depicted in Figure˜6(c).
@@
Figure 6: The values of $\operatorname{FP}(\mathbf{a}),\operatorname{FN}(\mathbf{a})$ and $\operatorname{err}(\mathbf{a})$ on the half-open chambers of the data arrangement for the $1$-dimensional dataset from Section˜A.2. The diagonal lines in all three images show the function $\lambda\mapsto\mathbf{a}^{*}(\lambda)$ defined in Section˜3.2.
Figure˜7 shows the level sets of the log-likelihood loss function for the same example, for two choices of the rate parameter $\lambda$. In accordance to Section˜3.2, the plots illustrate that the superlevel sets of these functions are convex. Moreover, the rate parameter has an influence on the (unique) maximizer of these functions, but they lie on a line $\lambda\mapsto\mathbf{a}^{*}(\lambda)$, as shown in Theorem˜3.11.
The same line is drawn in Figures˜6(a) and 6(b), illustrating that the numbers of false positives and false negative are monotone increasing and decreasing, respectively (cf. Section˜3.2). In Figure˜6(c), the line crosses regions where the $0/1$-loss has values $3,2,3,2,3,2,3,4$, respectively. This shows that the $0/1$-loss is not “convex” in the sense that the sequence of values of the $0/1$-loss along this line is unimodal, but goes up and down repeatedly.
@@
Figure 7: Level sets of log-likelihood loss functions for different choices of $\lambda$ on the dataset from Section˜A.2. The black dot depicts the minimum $\mathbf{a}^{*}(\lambda)$, lying on the line of maxima when varying the choice of $\lambda$.
A.3 Examples of the geometry of the generalized parameter space (Section˜4)
@@
Figure 8: The data points from Section˜A.3, with fixed $\mathbf{a}$ and two perfect classifiers.
Example A.4 (Star arrangement).
Consider the $2$-dimensional labeled dataset $X$ consisting of $3$ distinct points
$(\mathbf{x}^{(1)},y^{(1)})=(1,1,0),\quad(\mathbf{x}^{(2)},y^{(2)})=(2,2,1),\quad(\mathbf{x}^{(3)},y^{(3)})=(3,3,0),$
and let $\Delta$ be the $2$-dimensional Coxeter fan of type B (cf. Section˜A.1), which has $8$ rays. We fix the star through values
$\mathbf{a}=\left(\frac{1}{3},3,\frac{1}{3},3,\frac{1}{3},3,\frac{1}{3},3\right).$
Figure˜8 shows two different translations $\operatorname{star}(\mathbf{a})+\mathbf{t}$ that are perfect classifiers. Figure˜9 shows the associated star arrangement in $\mathbb{R}^{2}$, and the sublevel sets of the translational 0/1-loss. In particular, in contrast to Theorem˜3.5, it can be observed that the $0^{\text{th}}$ level set $L(\operatorname{err}_{\mathbf{a}},0)$ consists of two full-dimensional connected components. The translated stars in Figure˜8 show one example from each of the connected components.
223333333333332222222222222222111111111100-2-2-1-100112233445566
Figure 9: The star arrangement from Section˜A.3, and the level sets of the translational 0/1-loss function.
Appendix B Experiments
In this section we collect illustrations of the results of our experiments described in Section 5.
(a) Algorithm 1
(b) Neural network.
(c) SVM with RBF kernel.
Figure 10: Decision boundaries for Algorithm 1, a neural network and SVM with RBF kernel.
(a) Regular fan on 88 rays.
(b) Fan refinement.
(c) Fan coarsening.
Figure 11: Left: Fan on $8$ rays (blue) and the supported star (green) used in the generation of the synthetic data; Center: Refinement of the fan with one extra ray, and an example of a supported star; Right: Coarsening of the fan with one ray removed, and an example of a supported star.
(a) Regular fan on 88 rays.
(b) Fan refinement.
(c) Fan coarsening.
Figure 12: Optimal starshaped decision boundaries for Algorithm 1, with the same underlying fan as the synthetic data (left), a refinement of the fan (center) and a coarsening of the fan (right). The underlying fans are depicted in Figure˜11.
Appendix C Proofs
C.1 Proof of Theorem˜3.1
@@ Appendix C Proofs
Then $f_{\mathbf{a}}^{\Delta}(\mathbf{v}_{i})=a_{i}\geq 1$ if and only if $\ell(\mathbf{v}_{i})=1$, and thus $c_{\mathbf{a}}(\mathbf{v}_{i})=\ell(\mathbf{v}_{i})$ as claimed. It follows that the set of classifiers shatters the set of generators and thus the VC dimension of $S^{\Delta}$ is at least $n$.
To see that the VC dimension is at most $n$, we assume that there is a set $\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n+1)}\in\mathbb{R}^{d}$ of $n+1$ points that can be shattered by $S^{\Delta}$. By construction, for all $1\leq i\leq n+1$, $c_{\mathbf{a}}(\mathbf{x}^{(i)})=0$ if and only if $\langle[\mathbf{x}^{(i)}]^{\Delta},\mathbf{a}\rangle\leq 1$. In particular, if $\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n+1)}$ can be shattered by $S^{\Delta}$ then $\mathbf{x}^{(1)}]^{\Delta},\ldots,[\mathbf{x}^{(n+1)}]^{\Delta}$ can be shattered by the set of half-spaces of the form $\{\mathbf{x}\in\mathbb{R}^{n}\colon\mathbf{a}^{T}\mathbf{x}\leq b,a_{1},\ldots,a_{n}\geq 0\}$. This is a subset of the set of halfspaces considered in Section˜C.1 below, and thus we obtain a contradiction by Section˜C.1. This completes the proof. ∎
Proposition C.1.
@@
Then the VC dimension of $\mathcal{H}$ equals $n$.
@@
We need to show that there is no set of $n+1$ that can be shattered by $\mathcal{H}$. Let $\mathbf{y}_{1},\ldots,\mathbf{y}_{n+1}$ be points in $\mathbf{R}^{n}$ and let $\tilde{\mathbf{y}}_{1},\ldots,\tilde{\mathbf{y}}_{n+1}$ be their projections on the first n1n-1 coordinates. Then $\tilde{\mathbf{y}}_{1},\ldots,\tilde{\mathbf{y}}_{n+1}$ are affinely dependent, that is, there exist $\mu_{1},\ldots,\mu_{n}$, not all equal to zero, such that $\sum_{i}\mu_{i}=0$ and $\sum_{i}\mu_{i}\tilde{\mathbf{y}}_{i}=\mathbf{0}$. Let
[…]
@@
Then both $\mathbf{w}_{+}$ and $\mathbf{w}_{-}$ agree with $\tilde{\mathbf{w}}$ on the first n1n-1 coordinates. W.l.o.g we may assume that $(w_{-})_{n}\leq(w_{+})_{n}$, that is, the last coordinate of $w_{-}$ is not bigger than the last coordinate of $\mathbf{w}_{+}$. Then we claim that there exists no half-space $\{\mathbf{x}\in\mathbb{R}^{n}\colon\mathbf{a}^{T}\mathbf{x}\leq b,a_{n}\geq 0\}$ in $\mathcal{H}$ such that $\mathbf{a}^{T}\mathbf{y}_{i}\leq b$ for all $i$ such that $\mu_{i}>0$ and $\mathbf{a}^{T}\mathbf{y}_{i}>b$ for all $i$ such that $\mu_{i}<0$. To see this, we assume to the contrary that such a hyperplane exists. We then have
@@
and similarly $\mathbf{a}^{T}\mathbf{w}_{-}>b$. In particular, $\mathbf{a}^{T}\mathbf{w}_{-}>\mathbf{a}^{T}\mathbf{w}_{+}$. Since $\mathbf{w}_{+}$ and $\mathbf{w}_{-}$ agree on the first n1n-1 coordinates it therefore follows that $a_{n}(w_{+})_{n}<a_{n}(w_{-})$ and thus $\mathbf{w}_{+}<\mathbf{w}_{-}$ since $a_{n}>0$, a contradiction. Thus, no hyperplane in $\mathcal{H}$ satisfies the claim and thus $\mathcal{H}$ does not shatter any set of $n+1$ points. The VC dimension is thus at most $n$. To see that it is fact equal to $n$ we observe that the set of unit vectors $\mathbf{e}_{1},\ldots,\mathbf{e}_{n}$ can be shattered. ∎
C.2 Proof of Theorem˜3.4
@@
From the proof of Theorem˜3.6 we see that the Hessian of $\mathcal{L}(\mathbf{a})$ is a negative linear combination of the rank-$1$ matrices $\left([\mathbf{x}^{(i)}]^{\Delta}\right)^{T}[\mathbf{x}^{(i)}]^{\Delta}$ for $\mathbf{x}^{(i)}\in X_{1}$, that is, $\operatorname{Hess}\mathcal{L}(\mathbf{a})=\sum_{i}\lambda_{i}\left([\mathbf{x}^{(i)}]^{\Delta}\right)^{T}[\mathbf{x}^{(i)}]^{\Delta}$ for some $\lambda_{i}<0$ where the sum is over all $i$ such that $\mathbf{x}^{(i)}\in X_{1}$. Now let $\mathbf{v}$ be an eigenvector of $\operatorname{Hess}\mathcal{L}(A)$ with eigenvalue $\mu$. Since $A_{X_{1}}$ has rank $n$, there exists an $\mathbf{x}^{(i_{0})}\in X_{1}$ with $\langle[\mathbf{x}^{(i_{0})}]^{T},\mathbf{v}\rangle\neq 0$. Then
@@
It follows that $\mu\neq 0$ and thus $\mu<0$. Thus, $\operatorname{Hess}\mathcal{L}(A)$ is negative definite and thus $\mathcal{L}(A)$ is strictly concave. Since the parameter space $\mathbb{R}_{>0}^{n}=\left\{\mathbf{a}:\mathbf{a}>0\right\}$ is convex, $\operatorname{Hess}\mathcal{L}(A)$ has a unique maximum on the extended positive orthant $(\mathbb{R}_{\geq 0}\cup\{\infty\})^{n}$. If furthermore the rank of $A_{X_{0}}$ is $n$ then for each $j\in\{1,\dots,n\}$ there exists an $x^{(i)}\in X_{0}$ such that $\mathbf{x}^{(i)}]_{j}>0$. Therefore, we see that $\mathcal{L}(\mathbf{a})\rightarrow$-$\infty$ whenever $a_{j}\rightarrow\infty$. Thus, the unique maximum must be attained in the closed positive orthant $(\mathbb{R}_{\geq 0})^{n}$. ∎
C.6 Proof of Theorem˜3.11
@@
From the proof of Theorem˜3.11 we see that for any $0<\lambda<\lambda^{\prime}$ holds $\mathbf{a}^{*}(\lambda^{\prime})=\lambda/\lambda^{\prime}\cdot\mathbf{a}^{*}(\lambda)<\mathbf{a}^{*}(\lambda)$. Since $\mathbf{a}^{*}(\lambda)\in\mathbb{R}_{>0}^{n}$ both claims follow from Theorem˜3.4. ∎
C.8 Proof of Section˜4.1
@@
We begin with the first statement. First, let $C=\operatorname{cone}(\mathbf{v}_{i_{1}},\dots,\mathbf{v}_{i_{d}})$ be a full-dimensional cone of $\Delta$ such that $\mathbf{t}$ is contained in the interior of the cone $\mathbf{x}^{(i)}-C$ of the fan $\mathbf{x}^{(i)}-\Delta$. Equivalently, $\mathbf{x}^{(i)}-\mathbf{t}\in C$. With $V_{C}=(\mathbf{v}_{i_{1}}\dots\mathbf{v}_{i_{d}})$, we can thus compute $\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta}=V_{C}^{-1}(\mathbf{x}^{(i)}-\mathbf{t})$, which is a linear function in $\mathbf{t}$. The statement for lower-dimensional cones follows by taking limits. The second statement follows from substitution of the variables $\mathbf{x}\mapsto-\mathbf{t}$ and $\mathbf{t}\mapsto\mathbf{x}^{(i)}$ in (3). ∎
C.9 Proof of Theorem˜4.3
@@
Section˜4.1 implies that
@@
By Section˜4.1, for any $i$, $\mathbf{t}\mapsto f_{\mathbf{a}}^{\Delta}(\mathbf{x}^{(i)}-\mathbf{t})=\langle\mathbf{a},[\mathbf{x}^{(i)}-\mathbf{t}]\rangle$ is linear in $\mathbf{t}$ on every maximal cell $F$ of the fan arrangement of $\mathbf{x}^{(i)}-\Delta$, $i=1,\ldots,m$. Let $g_{F}(\mathbf{t})=\langle\mathbf{a},[\mathbf{x}^{(i)}-\mathbf{t}]\rangle$ be this linear function. Then the Hessian of $\mathcal{L}_{\mathbf{a}}(\mathbf{t})$ is the sum of matrices of the form
@@
The first set equals $\mathbb{R}^{n}_{>0}\times(-C+\mathbf{x}^{(i)})$. Restricted to $(\mathbf{a},\mathbf{t})$ such that $\mathbf{t}\in\mathbf{x}^{(i)}-C$, the function $\mathbf{t}\mapsto[\mathbf{x}^{(i)}-\mathbf{t}]$ is a linear map by Section˜4.1, so the expression $f^{\Delta}_{\mathbf{a}}(\mathbf{x}^{(i)}-\mathbf{t})=\langle[\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta},\mathbf{a}\rangle$ is a quadratic polynomial in variables $t_{1},\dots,t_{d},a_{1},\dots,a_{n}$. Therefore, $\mathcal{S}^{0}(C,\mathbf{x}^{(i)})$ it the intersection of solutions to the linear inequalities defining the polyhedral cone $\mathbb{R}^{n}_{>0}\times(-C+\mathbf{x}^{(i)})$, and the quadratic inequality $\langle[\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta},\mathbf{a}\rangle\leq 1$. Similarly, $\mathcal{S}^{1}(C,\mathbf{x}^{(i)})$ is the intersection of $\mathbb{R}^{n}_{>0}\times(-C+\mathbf{x}^{(i)})$ with the set of solutions to the inequality $\langle[\mathbf{x}^{(i)}-\mathbf{t}]^{\Delta},\mathbf{a}\rangle>1$. ∎
C.12 Proof of Theorem˜4.7
@@
We consider the subdivision of $\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$ into sets
@@
where we range over all possible $b(C,i)\in\{0,1\}$ for all $C\in\Delta,i\in\{1,\dots,m\}$. By Section˜4.2, each $S^{b(C,i)}(C,\mathbf{x}^{(i)})$ is basic semialgebraic with defining polynomials of degree at most $2$, and hence the same holds for the above finite intersection. By construction, the extended $0/1$-loss is constant on each $\bigcap_{C\in\Delta}\ \bigcap_{i=1}^{m}\mathcal{S}^{b(C,i)}(C,\mathbf{x}^{(i)})$. Fix $k\in\mathbb{Z}_{\geq 0}$. Then the $k^{\text{th}}$ level set $L(\operatorname{err},k)$ of the extended $0/1$-loss is the (finite) union over all sets $\bigcap_{C\in\Delta}\ \bigcap_{i=1}^{m}\mathcal{S}^{b(C,i)}(C,\mathbf{x}^{(i)})$ on which the extended $0/1$-loss is equal to $k$. Thus, the level set is a semialgebraic set. Since sublevel sets are finite unions of level sets, the same holds for sublevel sets. ∎
C.13 Proof of Theorem˜4.8